OpenStudy (anonymous):

12.Factor: 12x^5y^7 – 8x^4y^9

OpenStudy (anonymous):

@Mathbreaker @johnweldon1993

OpenStudy (johnweldon1993):

12) \(\large 12x^5y^7 – 8x^4y^9\) Looks like both terms are divisible by 4 (12/4 = 3 and 8/4 = 2) Looks like there is at least x^4 in both terms...and at least y^7 in each term...so all those can be factored out... \[\large 4x^4y^7(3x - 2y^2)\] Make sense?

OpenStudy (anonymous):

yes

OpenStudy (johnweldon1993):

Now for 13) \[\huge \frac{12x^6 – 8x^5 + 4x }{4x}\] Well...this can be broken up into \[\large \frac{12x^6}{4x} - \frac{8x^5}{4x} + \frac{4x}{4x}\] So do all those calculations

OpenStudy (anonymous):

ok, hold on

OpenStudy (anonymous):

12x^6-8x^5+x^2

OpenStudy (johnweldon1993):

Not quite... So 12x^6/4x would be 3x^5 You can think about it like this... 12/4 = 3 x^6 / x would be taking away 1 x...so it would become x^5 and so on..